# Feel-Good Thompson Sampling for Contextual Bandits: a Markov Chain Monte Carlo Showdown

### v1

@@ 7 Conclusion
We present the first systematic evaluation of Feel-Good Thompson Sampling and its smoothed variant (SFG-TS) under both exact and approximate MCMC-derived posterior regimes. We run ablation studies on the impact of preconditioning, feel-good bonuses, and inverse temperature scaling. Our experiments on eight real-world and synthetic benchmarks indicate that high-fidelity posteriors combined with FGTS’ optimism bonus reduces regret in the linear and logistic setting, but that Neural bandits performs below par with vanilla TS. Finally, our studies point to specific instances where larger bonuses can cause a critical performance degeneracy in approximate posterior settings.
Limitations. Our experiments exhibit substantial run-to-run variability: on MNIST and Covertype, the cumulative regret spans an order of magnitude, even across seeds. This reflects the sensitivity of the samplers to learning-rate schedules, mini-batch noise, and other training hyperparameters. While we tuned these settings once per dataset using a modest grid, we do not perform an exhaustive search. Secondly, Luo and Bayati (2025) recently showed that certain geometry-aware variants of Thompson Sampling achieve the minimax optimal regret in linear contextual bandits. Our experiments place a large focus on evaluating the performance of various MCMC methods in practice; however, it would be interesting to rigorously benchmark the geometry-aware algorithms. Likewise, it would be interesting to also study and benchmark recent noise-adaptive variants of Thompson Sampling (Xu et al., 2023) with MCMC methods. Finally, it would be interesting to also benchmark the performance of MCMC Thompson sampling variants on generalized linear reward models (Li et al., 2017).
Societal Impacts. This work is empirical and foundational in nature. While it enables better understanding of the behavior of bandit algorithms in numerous synthetic and real-world benchmarks, it is not tied to any specific applications or deployments.
@@ 8 Acknowledgements
This work was supported by NSF Grant CCF 2338816. We express our thanks to Eric Mazumdar, Jan van den Brand, Jake Abernethy, and Kabeer Thockchom for insightful discussions.
@@ Appendix B Relevant Literature
Regret bounds in Linear Contextual Bandits. Linear contextual bandits have been extensively studied, particularly in relation to regret bounds achieved by policies based on the Upper Confidence Bound (UCB) principle, where frequentist regret bounds of $\widetilde{O}(d\sqrt{T})$ are attainable [Zhou et al., 2020]. Thompson Sampling (TS), which involves sampling from the posterior distribution of the reward function and selecting actions accordingly, demonstrates strong empirical performance and applies naturally to both Bayesian and frequentist frameworks. Russo and Van Roy Russo and Roy [2016] showed that the best-known Bayesian regret bound for TS matches the minimax optimal $\widetilde{O}(d\sqrt{T})$. However, the best frequentist regret bound proven to date for TS is $O(d\sqrt{dT})$ Agrawal and Goyal [2013] due primarily to inflated posterior variance of order $\widetilde{O}(d)$ Significant contributions to this area leverage concentration inequalities for self-normalized martingales, as seen in Bercu and Touati [2019], Abbasi-Yadkori et al. [2011], Agrawal and Goyal [2013], Bietti et al. [2021], Agarwal et al. [2021, 2014]. There is also a body of work that shows that in cases where structural assumptions are made on the linear contextual bandit problem, there are algorithms that achieve the optimal regret bound; for instance, Chu et al. [2011] studies the case with linear payoff functions and Reid et al. [2025] studies the case with budget constraints. Both works derive the optimal $O(d\sqrt{T})$ regret bounds.
@@
Beyond Standard Linear Settings. Expanding beyond linear setting, substantial progress has been made in generalized linear bandits [Filippi et al., 2010], where the true reward $r$ from arm $x\in\mathcal{X}_{t}$ at round $t$ is assumed to be from a generalized linear model (GLM). In a GLM, conditional on feature vector $x$, $r$ follows an exponential family distribution with mean $\mu(x^{\top}\theta^{*})$ where $\theta^{*}$ is an unknown weight parameter shared across all arms and $\mu(\cdot)$ is called the link function. Sivakumar et al. [2020] provides a geometric perspective on contextual bandits, introducing smoothness through Gaussian perturbations in adversarial contexts. Further, adaptive noise strategies for Thompson sampling have been analyzed rigorously, resulting in refined regret bounds [Xu et al., 2023].
@@ C.2 Proof of Convergence of Preconditioned SFGLMCTS
where $\mathcal{L}_{t}$ is the smoothed feel-good likelihood, and $\epsilon_{t,k}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is an isotropic Gaussian vector. Note that this is identical to the preconditioned LMCTS update by setting the feel-good parameter $\lambda$ in $\mathcal{L}_{t}$ to 00. The $\theta$-update in Equation 5 can be viewed as the Euler-Maruyama discretization of the preconditioned Langevin dynamics from physics given by the stochastic differential equation
@@
As (I) =absent=-= - (II), the density proportional to $e^{-\beta_{t}\mathcal{L}_{t}(\theta)}$ is the stationary distribution for the SDE. Since $L_{t}$ and $\beta_{t}$ are measurable, Lemma 5.2.1 in Øksendal [2003] (or Theorem B.3.1 in Bakry et al. [2013]) gives that the SDE admits a unique stationary distribution, which proves the lemma.∎
@@ C.4 Analysis of Hamiltonian Monte Carlo Feel-Good Thompson Sampling
Under certain conditions on $-\nabla\mathcal{L}_{t}(\theta)$, the distribution of the Markov chain in the inner loop of Algorithm 3 converges to a unique stationary distribution $\pi(\mathrm{d}\theta)\propto e^{-\mathcal{L}_{t}(\theta)}\mathrm{d}\theta$ (see Lemma C.6). We provide details on sampling $p^{\prime}$ using numerical leapfrog integrators in Algorithm 4, and show in Theorem C.4 that HMC-FG-TS in linear contextual bandits satisfies a regret of $\widetilde{O}(d\sqrt{T})$.
@@
C.5 Proof of Convergence of SFGHMCTS
@@
If $g:N\to M$ is a $C^{k}$-function (i.e., $k$ times continuously differentiable) for $k\geq\max\{n-m+1,1\}$ and if $A_{r}\subseteq N$ is the set of points $x\in N$ such that $dg_{x}$ has rank $<r$, then the $r$-dimensional Hausdorff measure of $g(A_{r})$ is zero.
@@
Since the kernel of the HMC Markov chain is ergodic, we have that there exists a unique stationary distribution. The reversibility of HMC implies that $\exp(-H(\theta,v))$ is the (unique) stationary distribution of the joint Markov chain on $(\theta,v)$. Therefore,
@@ Appendix D Algorithm Descriptions
SVRGLMCTS $\eta=1$, batch size $=64$.
@@
PHMCTS $\eta=1$, leapfrog-step $=10$.
FGHMCTS $\lambda=0.5$, $\eta=1$, $b=1000$, leapfrog-step $=10$.
PFGHMCTS $\lambda=0.5$, $\eta=1$, $b=1000$, $\lambda_{\text{reg}}=1$, leapfrog-step $=10$.
SFGHMCTS $\lambda=0.5$, $\eta=1$, $b=1000$, $s=10$, leapfrog-step $=10$.
PSFGHMCTS $\lambda=0.5$, $\eta=1$, $b=1000$, $s=10$, $\lambda_{\text{reg}}=1$, leapfrog-step $=10$.
@@
SVRGLMCTS $\eta=1$, batch size $=64$.
@@
Table 10: Detailed Description of the Algorithms in the Neural Experiments: Covertype.
@@
NeuralTS $\ell_{s}$ loss, ReLU, $\nu=0.01$, $\lambda_{\text{reg}}=0.001$.
FG-NeuralTS $\ell_{s}$ loss, ReLU, $\nu=0.01$, $\lambda_{\text{reg}}=0.001$, $\lambda=1.0$, $b=1.0$.
SFG-NeuralTS $\ell_{s}$ loss, ReLU, $\nu=0.01$, $\lambda_{\text{reg}}=0.001$, $\lambda=1.0$, $b=1.0$, $s=10$.
Table 11: Detailed Description of the Algorithms in the Neural Experiments: Magic.
@@
NeuralTS $\ell_{2}$ loss, ReLU, $\nu=0.00001$.
FG-NeuralTS $\ell_{2}$ loss, ReLU, $\nu=0.00001$, $\lambda=1$, $b=1$.
SFG-NeuralTS $\ell_{2}$ loss, ReLU, $\nu=0.00001$, $\lambda=1$, $b=1$, $s=10$.
@@ E.1 Linear Contextual Bandits
Low-dimensional setting. For the $d=20$ setting, Tables 17 and 18 provide our experiments over linear contextual bandits for $\beta=10^{3}$ and $\beta=1$ (respectively), where we modify the feel-good parameter $\lambda$ in the loss-likelihood for values $\lambda\in\{0.01,0.1,0.5,1.0\}$. The corresponding vanilla Thompson sampling cumulative regret values are equivalent to setting $\lambda=0$. For these values, we direct the reader to Table 1. These experiments strongly indicate that setting $\lambda=0.01$ outperforms the case where $\lambda=0$ (as well as other choices of $\lambda$). Especially for SFGMALATS, which enjoys optimal regret in this setting, the regret is considerably smaller than that of LMCTS and FGLMCTS.
@@
SFGMALATS 70.3 $\pm$ 43.5 669.1 $\pm$ 1839.6 366 $\pm$ 183.6 758.7 $\pm$ 303.5
High-dimensional setting. For the $d=40$ setting, Tables 19 and 20 provide our experiments over linear contextual bandits for $\beta=10^{3}$ and $\beta=1$ (respectively), where we modify the feel-good parameter $\lambda$ in the loss-likelihood for values $\lambda\in\{0.01,0.1,0.5,1.0\}$. As before, we see that setting the feel-good parameter $\lambda$ to be $0.01$ outperforms the case where $\lambda=0$ (as well as other choices of $\lambda)$. As before, this particularly carries forward for FGMALATS and SFGMALATS which experience the lowest cumulative regrets.
@@
SFGLMCTS $4499.2\pm 118.1$ $4486.4\pm 135.7$ $4406.7\pm 139.6$ $4395.4\pm 166.3$
SFGMALATS $100.1\pm 19.2$ $110.6\pm 12.7$ $130.1\pm 16.7$ $425.7\pm 155.3$

### v3

@@ 7 Conclusion
We present the first systematic evaluation of FGTS and its smoothed variants under exact and approximate MCMC-derived posterior regimes. Our ablations analyze the effects of preconditioning, feel-good bonuses, and temperature scaling. Across 14 real-world and synthetic datasets, we find that high-fidelity posteriors combined with FGTS’s optimism reduce regret in linear and logistic settings, whereas neural bandits underperform compared to vanilla TS. We also identify cases where larger bonuses cause severe performance degradation under approximate posteriors.
Limitations and Future Work. Our experiments reflect the sensitivity of the samplers to learning-rate schedules, mini-batch noise, and other hyperparameters. While we used a modest grid search for tuning, our study was not exhaustive. Recent work (Luo and Bayati, 2025) shows that geometry-aware variants of TS can achieve minimax-optimal regret in linear bandits. Benchmarking such algorithms, along withs noise-adaptive TS variants (Xu et al., 2023) and MCMC-based methods on generalized linear reward models (Li et al., 2017), would be a promising direction for future work.
Societal Impacts. This work is empirical in nature. While it enables better understanding of the behavior of bandit algorithms in numerous synthetic and real-world benchmarks, it is not tied to any specific applications or deployments.
@@ 8 Acknowledgements
This work was supported by NSF Grant CCF 2338816. We express our thanks to Eric Mazumdar, Jan van den Brand, Jake Abernethy, Haque Ishfaq, Yunbum Kook, Mirabel Reid, and Kabeer Thockchom for insightful discussions.
@@ Appendix B Relevant Literature
Regret bounds in Linear Contextual Bandits. Linear contextual bandits have been extensively studied, particularly in relation to regret bounds achieved by policies based on the Upper Confidence Bound (UCB) principle, where frequentist regret bounds of $\widetilde{O}(d\sqrt{T})$ are attainable [Zhou et al., 2020]. Thompson Sampling (TS), which involves sampling from the posterior distribution of the reward function and selecting actions accordingly, demonstrates strong empirical performance and applies naturally to both Bayesian and frequentist frameworks. Russo and Roy [2016] showed that the best-known Bayesian regret bound for TS matches the minimax optimal $\widetilde{O}(d\sqrt{T})$. However, the best frequentist regret bound proven to date for TS is $O(d\sqrt{dT})$ [Agrawal and Goyal, 2013] due primarily to inflated posterior variance of order $\widetilde{O}(d)$ Significant contributions to this area leverage concentration inequalities for self-normalized martingales, as seen in Bercu and Touati [2019], Abbasi-Yadkori et al. [2011], Agrawal and Goyal [2013], Bietti et al. [2021], Agarwal et al. [2021, 2014]. There is also a body of work that shows that in cases where structural assumptions are made on the linear contextual bandit problem, there are algorithms that achieve the optimal regret bound; for instance, Chu et al. [2011] studies the case with linear payoff functions and Reid et al. [2025] studies the case with budget constraints. Both works derive the optimal $O(d\sqrt{T})$ regret bounds.
@@
Beyond Standard Linear Settings. Expanding beyond linear setting, substantial progress has been made in generalized linear bandits [Filippi et al., 2010], where the true reward $r$ from arm $x\in\mathcal{X}_{t}$ at round $t$ is assumed to be from a generalized linear model (GLM). In a GLM, conditional on feature vector $x$, $r$ follows an exponential family distribution with mean $\mu(x^{\top}\theta^{*})$ where $\theta^{*}$ is an unknown weight parameter shared across all arms and $\mu(\cdot)$ is called the link function. Sivakumar et al. [2020] provides a geometric perspective on contextual bandits, introducing smoothness through Gaussian perturbations in adversarial contexts. Further, adaptive noise strategies for Thompson sampling have been analyzed rigorously, resulting in refined regret bounds [Ishfaq et al., 2024b, a, Xu et al., 2023].
@@ C.2 Proof of Convergence of Preconditioned SFGLMCTS
where $\mathcal{L}_{t}$ is the smoothed feel-good likelihood, and $\epsilon_{t,k}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is an isotropic Gaussian vector. Note that this is identical to the preconditioned LMCTS update by setting the feel-good parameter $\lambda$ in $\mathcal{L}_{t}$ to 0. The $\theta$-update in Equation 5 can be viewed as the Euler-Maruyama discretization of the preconditioned Langevin dynamics from physics given by the stochastic differential equation
@@
As (I) ==- (II), the density proportional to $e^{-\beta_{t}\mathcal{L}_{t}(\theta)}$ is the stationary distribution for the SDE. Since $L_{t}$ and $\beta_{t}$ are measurable, Lemma 5.2.1 in Øksendal [2003] (or Theorem B.3.1 in Bakry et al. [2013]) gives that the SDE admits a unique stationary distribution, which proves the lemma.∎
@@ C.4 Analysis of Hamiltonian Monte Carlo Feel-Good Thompson Sampling
Under certain conditions on $\nabla\mathcal{L}_{t}(\theta)$, the distribution of the Markov chain in the inner loop of Algorithm 3 converges to a unique stationary distribution $\pi(\mathrm{d}\theta)\propto e^{-\mathcal{L}_{t}(\theta)}\mathrm{d}\theta$ (see Lemma C.6). We provide details on sampling $p^{\prime}$ using numerical leapfrog integrators in Algorithm 4, and show in Theorem C.4 that HMC-FG-TS in linear contextual bandits satisfies a regret of $\widetilde{O}(d\sqrt{T})$.
@@
C.5 Proof of Convergence of Smoothed Feel-Good HMCTS
@@
If $g:N\to M$ is a $C^{k}$-function (i.e., $k$ times continuously differentiable) for $k\geq\max\{n-m+1,1\}$ and if $A_{r}\subseteq N$ is the set of points $x\in N$ such that $dg_{x}$ has rank <r<r, then the $r$-dimensional Hausdorff measure of $g(A_{r})$ is zero.
@@
Since the kernel of the HMC Markov chain is ergodic, we have that there exists a unique stationary distribution [Anand and Umans, 2023]. The reversibility of HMC implies that $\exp(-H(\theta,v))$ is the (unique) stationary distribution of the joint Markov chain on $(\theta,v)$. Therefore,
@@ Appendix D Algorithm Descriptions
SVRGLMCTS $\eta=1$, batch size =64=64.
@@
PHMCTS $\eta=1$, leapfrog-step =10=10.
FGHMCTS $\lambda=0.5$, $\eta=1$, $b=1000$, leapfrog-step =10=10.
PFGHMCTS $\lambda=0.5$, $\eta=1$, $b=1000$, $\lambda_{\text{reg}}=1$, leapfrog-step =10=10.
SFGHMCTS $\lambda=0.5$, $\eta=1$, $b=1000$, $s=10$, leapfrog-step =10=10.
PSFGHMCTS $\lambda=0.5$, $\eta=1$, $b=1000$, $s=10$, $\lambda_{\text{reg}}=1$, leapfrog-step =10=10.
@@
SVRGLMCTS $\eta=1$, batch size =64=64.
@@
Table 10: Detailed Description of the Algorithms in the Neural Experiments: Covertype ($\ell_{2}$ loss).
@@
NeuralTS ReLU, $\nu=0.01$, $\lambda_{\text{reg}}=0.001$.
FG-NeuralTS ReLU, $\nu=0.01$, $\lambda_{\text{reg}}=0.001$, $\lambda=1.0$, $b=1.0$.
SFG-NeuralTS ReLU, $\nu=0.01$, $\lambda_{\text{reg}}=0.001$, $\lambda=1.0$, $b=1.0$, $s=10$.
Table 11: Detailed Description of the Algorithms in the Neural Experiments: Magic ($\ell_{2}$ loss).
@@
NeuralTS ReLU, $\nu=0.00001$.
FG-NeuralTS ReLU, $\nu=0.00001$, $\lambda=1$, $b=1$.
SFG-NeuralTS ReLU, $\nu=0.00001$, $\lambda=1$, $b=1$, $s=10$.
@@
Table 17: Detailed Description of the Algorithms in the Neural Experiments: RestaurantRatings ($\ell_{2}$ loss).
Algorithm Description
Neural-$\epsilon$-Greedy $\epsilon=0.01$.
LMCTS $\beta^{-1}=0.0001$, LeakyReLU.
FGLMCTS $\beta^{-1}=0.0001$, LeakyReLU, $\lambda=0.01,b=1.0$.
SFGLMCTS $\beta^{-1}=0.0001$, LeakyReLU, $\lambda=0.01,b=1.0,s=10$.
NeuralUCB $\lambda_{\text{reg}}=0.01$.
NeuralTS ReLU, $\nu=0.00001$, $\lambda_{\text{reg}}=0.01$.
FG-NeuralTS ReLU, $\nu=0.00001$, $\lambda_{\text{reg}}=0.01$, $\lambda=0.01$, $b=1.0$.
SFG-NeuralTS ReLU, $\nu=0.00001$, $\lambda_{\text{reg}}=0.01$, $\lambda=0.01$, $b=1.0$, $s=10$.
Table 18: Detailed Description of the Algorithms in the Neural Experiments: CIFAR-10 ($\ell_{2}$ loss).
Algorithm Description
Neural-$\epsilon$-Greedy $\epsilon=0.01$.
LMCTS $\beta^{-1}=0.000001$, LeakyReLU.
FGLMCTS $\beta^{-1}=0.000001$, LeakyReLU, $\lambda=0.01,b=1.0$.
SFGLMCTS $\beta^{-1}=0.000001$, LeakyReLU, $\lambda=0.01,b=1.0,s=10$.
NeuralUCB $\lambda_{\text{reg}}=0.01$.
NeuralTS ReLU, $\nu=0.00001$, $\lambda_{\text{reg}}=0.01$.
FG-NeuralTS ReLU, $\nu=0.00001$, $\lambda_{\text{reg}}=0.01$, $\lambda=0.01$, $b=1.0$.
SFG-NeuralTS ReLU, $\nu=0.00001$, $\lambda_{\text{reg}}=0.01$, $\lambda=0.01$, $b=1.0$, $s=10$.
@@ E.1 Linear Contextual Bandits
Low-dimensional setting. For the $d=20$ setting, Tables 19 and 20 provide our experiments over linear contextual bandits for $\beta=10^{3}$ and $\beta=1$ (respectively), where we modify the feel-good parameter $\lambda$ in the loss-likelihood for values $\lambda\in\{0.01,0.1,0.5,1.0\}$. The corresponding vanilla Thompson sampling cumulative regret values are equivalent to setting $\lambda=0$. For these values, we direct the reader to Table 1. These experiments strongly indicate that setting $\lambda=0.01$ outperforms the case where $\lambda=0$ (as well as other choices of $\lambda$). Especially for SFGMALATS, which enjoys optimal regret in this setting, the regret is considerably smaller than that of LMCTS and FGLMCTS.
@@
SFGMALATS 70.3 $\pm$ 43.5 669.1 $\pm$ 1839.6 366.0 $\pm$ 183.6 758.7 $\pm$ 303.5
High-dimensional setting. For the $d=40$ setting, Tables 21 and 22 provide our experiments over linear contextual bandits for $\beta=10^{3}$ and $\beta=1$ (respectively), where we modify the feel-good parameter $\lambda$ in the loss-likelihood for values $\lambda\in\{0.01,0.1,0.5,1.0\}$. As before, we see that setting the feel-good parameter $\lambda$ to be $0.01$ outperforms the case where $\lambda=0$ (as well as other choices of $\lambda)$. As before, this particularly carries forward for FGMALATS and SFGMALATS which experience the lowest cumulative regrets.
@@
SFGLMCTS $499.2\pm 118.1 44$$86.4\pm 135.7 44$$406.7\pm 139.6$ $4395.4\pm 166.3$
SFGMALATS $00.1\pm 19.2 11$$10.6\pm 12.7$ 130.1 $\pm$ 16.7 425.7 $\pm$ 155.3
E.1.1 Posterior Analysis
A central theme of our work is that given MCMC-based TS variants compute an approximate posterior $\widetilde{\pi}_{t}$ through sampling procedures, we wish to measure how faithfully these MCMC approximations capture the true Bayesian posterior? To disentangle these factors, following Riquelme et al. [2018], we conduct a controlled posterior quality analysis where all algorithms observe identical data, allowing us to isolate approximation quality from exploration strategy. Our analysis is motivated by the following observation: if $\widetilde{\pi}_{t}$ is consistently close to the true posterior $\pi_{t}$ across timesteps and arms, then the MCMC approximation preserves the Bayesian reasoning that underpins TS’ theoretical properties. Conversely, significant divergence between $\widetilde{\pi}_{t}$ and $\pi_{t}$ suggests that the algorithm’s beliefs are systematically biased, potentially undermining exploration-exploitation balance.
We consider a linear contextual bandit with $K=6$ arms in $d=20$ dimensions, following the model proposed by Russo and Roy [2016]. For each arm $i\in\{1,\dots,K\}$, the reward function is:
$r_{i,t}=X_{t}^{\top}\beta_{i}+\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$
where $X_{t}\in\mathbb{R}^{d}$ is the context vector and $\beta_{i}\in\mathbb{R}^{d}$ is the arm-specific parameter. We set the prior $\beta_{i}\sim\mathcal{N}(0,\lambda^{-1}\mathbf{I}d)$ with $\lambda=1.0$ and observation noise $\sigma=0.5$.
To ensure a fair comparison across algorithms, we generate a fixed dataset that all algorithms observe. We first sample true parameters $\beta_{i}\sim\mathcal{N}(0,\mathbf{I}_{d})$ for each arm $i$, representing the ground truth (which is unknown to the algorithms). To generate the contexts, we form $T=2000$ context vectors with planted correlation structures to induce non-isotropic posteriors. We begin with $X_{t}^{(0)}\sim\mathcal{N}(0,\mathbf{I}_{d})$ and apply transformations $X_{t}^{(1)}\leftarrow 1.7\cdot X_{t}^{(1)}$ and $X_{t}^{(2)}\leftarrow 0.55\cdot X_{t}^{(2)}+0.6\cdot X_{t}^{(1)}$. This creates elliptical posteriors as seen in Figure 2. Next, to generate the rewards, for each time step $t$, each algorithm chooses action $a_{t}$ and observes reward $r_{t}=X_{t}^{\top}\beta_{a_{t}}+\epsilon_{t}$ where $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$.
Computing the true posterior. For the linear-Gaussian model, the posterior distribution admits a closed-form solution. Let $D_{t}^{(i)}=\{(X_{s},r_{s}):a_{s}=i,s\leq t\}$ denote the data collected for arm $i$ up to time $t$, with $n_{t}^{(i)}=|D_{t}^{(i)}|$ observations. The posterior for $\beta_{i}$ is Gaussian with
$\Lambda_{\text{post}}^{(i)}=\lambda I_{d}+\frac{\eta}{\sigma^{2}}\sum_{(X,r)\in D_{t}^{(i)}}XX^{\top},$ (9)
$\Sigma_{\text{post}}^{(i)}=(\Lambda_{\text{post}}^{(i)})^{-1},$ (10)
$\mu_{\text{post}}^{(i)}=\Sigma_{\text{post}}^{(i)}\cdot\frac{\eta}{\sigma^{2}}\sum_{(X,r)\in D_{t}^{(i)}}Xr,$ (11)
where $\eta$ is the inverse temperature parameter. This analytical posterior for each $i$ is $(\pi^{\star}_{t})^{(i)}\sim\mathcal{N}(\mu_{post}^{(i)},\Sigma_{post}^{(i)})$, which serves as our ground truth for comparison. The linear bandit problem with block-diagonal feature map $\phi(X,a)$ results in independent posteriors for each arm. When arm $j\neq i$ is played, the posterior $\beta_{i}$ remains unchanged. This allows us to analyze each arm’s posterior separately.
For each algorithm, we extract posterior samples at the final timestep $T$. LinTS maintains the exact posterior analytically via Sherman-Morrison updates. Therefore, we directly sample from $\mathcal{N}(\mu_{\text{LinTS}},\Sigma_{\text{LinTS}})$ to obtain 1500 posterior samples. However, for the MCMC variants such as LMCTS, MALATS, and preconditioned variants, we extract samples by repeatedly invoking the underlying MCMC sampling algorithm. Each call performs $K=100$ MCMC iterations (Langevin steps, MALA steps, or leapfrog integrations for HMC), producing one sample from the approximate posterior $\widetilde{\pi}_{t}$. We provide 2D scatter plots of $(\beta_{1},\beta_{2})$ projections showing 1500 samples from $\pi^{\star}_{t}$ (green) and $\widetilde{\pi}_{t}$ (red). Overlapping clouds indicate good approximation, whereas separation reveals bias.
Figure 2: MCMC Sampling with the True Linear Posteriors. From top to bottom, we plot the true posteriors in green, and the sampled posteriors of Feel-Good LMCTS, Linear TS, LMCTS, MALATS, and PLMCTS in blue.
Based on Figure 2, we see that an algorithm may exhibit high cumulative regret due to under-exploration or due to poor approximation. Posterior analysis disambiguates these failure modes. For instance, PLMCTS’ approximation of the true posterior is most egregious and displays a bias in its approximation; conversely, MALATS’ and LinTS’ posterior approximation are tight and match the true posterior. While LMCTS and FGLMCTS appear to be unbiased estimators of the true posterior, their variances can be much larger in under-explored arms.